2y/5 -3x/4 = 1/2 find y given that x = 2
Answer by amoresroy(361)

Substitute the values of x in the equation.
2y/5 - 3(2)/4 = 1/2
2y/5 - 6/4 = 1/2
2y/5 = 1/2 + 6/4
2y/5 = 2
2y = 5*2
2y = 10
y = 5